# Advancing Embodied Intelligence in Robotic-Assisted Endovascular Procedures: A Systematic Review of AI Solutions

### v1

@@ Abstract
Index Terms:
robotic-assisted endovascular procedures, medical image analysis, reinforcement learning
I Introduction
@@
The structure of this review is presented in Figure 1(d). Section II explores intelligent perception techniques in endovascular procedures, emphasizing recent progress in vascular segmentation and intervention device tracking. These advancements are essential for improving the precision of such interventions. Section III explores the integration of machine learning in robotic-assisted endovascular procedures, providing an in-depth analysis of its role in advancing intelligence and refining control for these complex operations. Section IV assesses the major challenges facing the field, encompassing the limited availability of high-quality coronary imaging data, the lack of a standardized evaluation framework, constraints on the scalability of perception and control methodologies, ethical and legal implications, and the complexities involved in incorporating these technologies into clinical workflows.
@@
II Intelligent Perception Techniques in Endovascular Procedures
@@
II-A Vascular Segmentation Approaches
@@
II-A1 Supervised Learning Methods
@@
II-A2 Weakly-supervised Learning
To address the challenge posed by limited labeled data, researchers have explored weakly-supervised strategies, particularly through the generation of pseudo-labels. These approaches aim to leverage vast amounts of unlabeled data while minimizing the reliance on costly manual annotations. Zhang et al. developed a weakly supervised framework that utilizes noisy labels and self-paced learning, achieving a Dice score of 0.821 [41]. Their method incorporates curriculum learning and progressive refinement to effectively manage noisy labels. Ma et al. introduced a self-supervised approach that employs adversarial learning without the need for manual labels, reporting a Dice score of 0.557 on the XCAD dataset [20]. Their method leverages label-free training and consistency regularization to learn representative features from unlabeled data. More recently, Wu et al. proposed DeNVeR, an unsupervised deformable neural representation method of video segmentation, demonstrating the potential of label-free learning [42]. This approach achieved a Dice score of 0.733 across multiple datasets (XCAD, CADICA, and XACV) by leveraging deformable modeling and temporal consistency. Kim et al. introduced a contrastive diffusion-based learning approach, achieving Dice scores of 0.661, 0.700, and 0.621 on the XCAD, 134XCA, and 30XCA datasets respectively [35]. Their method utilizes label-free diffusion models and contrastive learning to extract meaningful features from unlabeled data.
II-A3 Advanced Clinical Applications
@@
II-A4 Available Public Datasets
The availability of large-scale, publicly accessible datasets for coronary vascular segmentation remains limited, posing challenges for reproducibility and standardized benchmarking. Due to clinical regulations and privacy concerns, most current research relies on private datasets, further constraining the accessibility of standardized resources for broader evaluation and comparison. However, several valuable public datasets have been developed, as summarized in Table I. The XCAD dataset [20], intended for weakly-supervised learning, comprises 1621 training frames and 126 test angiograms. DCA1 and CHUAC, utilized in multi-dataset validation studies [32], provide 134 expert-annotated X-ray coronary angiograms and 30 diverse angiography images, respectively. The ASOCA CTA dataset [48] contributes 60 CCTA cases with expert annotations, suitable for the development of 3D segmentation algorithms. CCA-200 [47] presents 200 CCTA images annotated by professional radiologists, while CADICA [42] offers over 6,000 invasive coronary angiography images, supporting unsupervised learning approaches.
TABLE I: Publicly Available Datasets for Coronary Vessel Segmentation
@@ II-A4 Available Public Datasets
II-B Intervention Device Perception
@@
TABLE II: Comprehensive Review of Data-Driven Methods for Endovascular Procedure Scene Perception
@@
7 Thuy et al. [61] Coarse-to-fine strategy for coronary vessel segmentation; Key Features: (1) UNet for the main coronary artery extraction (2) the second UNet for coronary artery-extracted segmentation Dice: 0.7640 Private
@@
14 Navab et al. [34] Weakly-supervised Learning for coronary vessel segmentation Contrastive diffusion-based learning; Key Features: (1) UNet with channel attention module (2) Centerline auxiliary supervision module Dice: 0.8528 Private
15 Zhang et al. [41] Weakly-supervised segmentation with self-paced learning; Key Features: (1) Noisy label handling (2) Curriculum learning (3) Progressive refinement Dice: 0.821 Private
TABLE III: Comprehensive Review of Data-Driven Methods for Endovascular Procedure Scene Perception (Continued)
@@
TABLE IV: Comprehensive Review of Data-Driven Methods for Endovascular Procedure Scene Perception (Continued)
@@
46 Du et al. [16] Guidewire endpoint localization via pixel-adjacent relation; Key Features: (1) UNet segmentation (2) Eight-neighborhood search mean pixel errors: 2.02 Private
@@
TABLE V: Comprehensive Review of Data-Driven Methods for Endovascular Procedure Scene Perception (Continued)
@@
II-B1 Guidewire Detection and Tracking
@@
Recent research has emphasized clinical integration and real-time performance. Zhang et al. [37] developed a system for real-time intra-operative tracking with low latency (Dice: 0.899), addressing the critical need for seamless integration into clinical workflows. Their subsequent work [80] introduced the Jigsaw transformer framework, incorporating background reverse attention and a novel training strategy to effectively handle multi-wire scenarios (F1: 0.899 for single wire, 0.846 for dual wire).
Ravigopal et al. [81] advanced the field with a real-time pose estimation system that combines continuum kinematic modeling with pose optimization, achieving an accuracy of 0.900. This approach underscores the potential for more comprehensive guidewire tracking beyond simple detection.
@@
II-B2 Catheter Detection and Tracking
@@
II-C Multi-modal Registration and Integration in Cardiovascular Imaging
@@
TABLE VI: Summary of Data-Driven Control Methods for Endovascular Intervention Robots
@@
TABLE VII: Summary of Data-Driven Control Methods for Endovascular Intervention Robots (continued)
No. Authors (Year) Task Category Method Performance Metrics
@@
III Learning-Based Control for Robotic-Assisted Endovascular Procedures
@@ IV Challenges
IV-A Underdevelopment of Transformer-Based Models and Foundational Vision Models for Endovascular Procedures
@@
IV-B Scarcity of High-Quality Coronary Imaging Data
@@
IV-C Lack of Unified Workflow Recognition for Endovascular Procedures
@@
IV-D Absence of a Standardized, Large-Scale Benchmark
Recent advancements in large-scale benchmarking, exemplified by the Behavior-1K dataset [149] developed by Fei-Fei Li’s team, have significantly enhanced performance evaluation in endovascular interventions (EIs) by capturing 1,000 diverse, real-world activities within controlled simulation environments. The work deemed "Best Paper at ICRA 2024" also extended robotic learning through the introduction of standardized, high-quality datasets and models that support reproducibility and generalization across a wide range of robotic systems [150]. In contrast, endovascular robotics remains constrained by the scarcity of large-scale, domain-specific benchmarks that encapsulate the intricacies of human vasculature and the clinical complexity encountered in real-world settings. This limitation hampers systematic comparisons of algorithmic performance, impedes meaningful evaluation under varying clinical conditions, and results in fragmented, non-generalizable findings.
@@ IV-D Absence of a Standardized, Large-Scale Benchmark
IV-E Scalability of Perception and Learning-Based Control
@@
To address these scalability challenges, future research must concentrate on developing advanced perception algorithms capable of maintaining high accuracy and robustness across a wider range of anatomical variations and pathological conditions. One promising direction involves the integration of multi-modal imaging techniques, including ultrasound and DSA [154], combined with advanced deep learning algorithms, to enhance the system’s capacity to adapt to real-time physiological changes. Additionally, transfer learning and domain adaptation approaches may aid in overcoming the challenge of generalizing control algorithms across different patient datasets. These strategies have the potential to substantially improve the flexibility and clinical applicability of endovascular robotic systems, ensuring their reliable performance across diverse patient populations, and ultimately leading to improved procedural outcomes and patient safety.
IV-F Ethical and Legal Concerns
@@
Decision-making autonomy represents another critical issue. While AI-powered robotic systems can enhance precision, their "closed-box" nature, with opaque decision-making processes, raises concerns about clinician trust [157, 158]. Trustworthiness, defined as the reliability and predictability of a system, is crucial to ensure clinicians have confidence in the technology’s performance. Furthermore, XAI is essential for transparency, enabling clinicians to understand the rationale behind the robot’s decisions. This is particularly important in high-risk procedures, in which clinicians may need to intervene if the system’s decision-making process is unclear [159].
@@
IV-G Integration with Existing Clinical Workflows
Beyond the technical challenges, the integration of robotic systems into existing clinical workflows poses a critical barrier to their widespread adoption. Although the potential of endovascular robotics to enhance precision, decrease procedural variability, and improve patient outcomes is evident, the transition from traditional manual methods to fully automated or semi-automated systems presents substantial challenges beyond purely technological aspects. A pivotal issue is the reluctance of clinical staff to embrace new systems, particularly when they are perceived as "black boxes" - autonomous systems with decision-making processes that are difficult to interpret or explain [158]. In endovascular procedures, in which precision and safety are paramount, the inability to fully comprehend how a robotic system arrives at its decisions can elicit significant levels of anxiety among medical professionals. Clinicians require assurances that these systems can manage complex and unpredictable situations in a manner akin to experienced human practitioners, and that they would operate predictably in high-pressure clinical environments [163].

### v2

@@ Abstract
1 Introduction
@@
The structure of this review is presented in Figure 1(d). Section 2 explores intelligent perception techniques in endovascular procedures, emphasizing recent progress in vascular segmentation and intervention device tracking. These advancements are essential for improving the precision of such interventions. Section 3 explores the integration of machine learning in robotic-assisted endovascular procedures, providing an in-depth analysis of its role in advancing intelligence and refining control for these complex operations. Section 4 assesses the major challenges facing the field, encompassing the limited availability of high-quality coronary imaging data, the lack of a standardized evaluation framework, constraints on the scalability of perception and control methodologies, ethical and legal implications, and the complexities involved in incorporating these technologies into clinical workflows.
@@
2 Intelligent Perception Techniques in Endovascular Procedures
@@
2.1 Vascular Segmentation Approaches
@@
2.1.1 Supervised Learning Methods
@@
2.1.2 Weakly-supervised Learning
To address the challenge posed by limited labeled data, researchers have explored weakly-supervised strategies, particularly through the generation of pseudo-labels. These approaches aim to leverage vast amounts of unlabeled data while minimizing the reliance on costly manual annotations. Zhang et al. developed a weakly supervised framework that utilizes noisy labels and self-paced learning, achieving a Dice score of 0.821 [41]. Their method incorporates curriculum learning and progressive refinement to effectively manage noisy labels. Ma et al. introduced a self-supervised approach that employs adversarial learning without the need for manual labels, reporting a Dice score of 0.557 on the XCAD dataset [20]. Their method leverages label-free training and consistency regularization to learn representative features from unlabeled data.
More recently, Wu et al. proposed DeNVeR, an unsupervised deformable neural representation method of video segmentation, demonstrating the potential of label-free learning [42]. This approach achieved a Dice score of 0.733 across multiple datasets (XCAD, CADICA, and XACV) by leveraging deformable modeling and temporal consistency. Kim et al. introduced a contrastive diffusion-based learning approach, achieving Dice scores of 0.661, 0.700, and 0.621 on the XCAD, 134XCA, and 30XCA datasets respectively [35]. Their method utilizes label-free diffusion models and contrastive learning to extract meaningful features from unlabeled data.
2.1.3 Advanced Clinical Applications
@@
2.1.4 Available Public Datasets
The availability of large-scale, publicly accessible datasets for coronary vascular segmentation remains limited, posing challenges for reproducibility and standardized benchmarking. Due to clinical regulations and privacy concerns, most current research relies on private datasets, further constraining the accessibility of standardized resources for broader evaluation and comparison. However, several valuable public datasets have been developed, as summarized in Table 1. The XCAD dataset [20], intended for weakly-supervised learning, comprises 1621 training frames and 126 test angiograms. DCA1 and CHUAC, utilized in multi-dataset validation studies [32], provide 134 expert-annotated X-ray coronary angiograms and 30 diverse angiography images, respectively. The ASOCA CTA dataset [48] contributes 60 CCTA cases with expert annotations, suitable for the development of 3D segmentation algorithms. CCA-200 [47] presents 200 CCTA images annotated by professional radiologists, while CADICA [42] offers over 6,000 invasive coronary angiography images, supporting unsupervised learning approaches.
Table 1: Publicly Available Datasets for Coronary Vessel Segmentation
@@ 2.1.4 Available Public Datasets
2.2 Intervention Device Perception
@@
Table 2: Comprehensive Review of Data-Driven Methods for Endovascular Procedure Scene Perception
@@
7 Thuy et al. [61] Coarse‐to‐fine strategy for coronary vessel segmentation; Key Features: (1) UNet for the main coronary artery extraction (2) the second UNet for coronary artery‐extracted segmentation Dice: 0.7640 Private
Table 3: Comprehensive Review of Data-Driven Methods for Endovascular Procedure Scene Perception (Continued)
@@
14 Navab et al. [34] Contrastive diffusion-based learning; Key Features: (1) UNet with channel attention module (2) Centerline auxiliary supervision module Dice: 0.8528 Private
Table 4: Comprehensive Review of Data-Driven Methods for Endovascular Procedure Scene Perception (Continued)
@@
15 Zhang et al. [41] Weakly-supervised Learning for coronary vessel segmentation Weakly-supervised segmentation with self-paced learning; Key Features: (1) Noisy label handling (2) Curriculum learning (3) Progressive refinement Dice: 0.821 Private
@@
Table 5: Comprehensive Review of Data-Driven Methods for Endovascular Procedure Scene Perception (Continued)
@@
Table 6: Comprehensive Review of Data-Driven Methods for Endovascular Procedure Scene Perception (Continued)
@@
Table 7: Comprehensive Review of Data-Driven Methods for Endovascular Procedure Scene Perception (Continued)
@@
Table 8: Comprehensive Review of Data-Driven Methods for Endovascular Procedure Scene Perception (Continued)
@@
46 Du et al. [16] Guidewire endpoint localization via pixel‐adjacent relation; Key Features: (1) UNet segmentation (2) Eight‐neighborhood search mean pixel errors: 2.02 Private
@@
Table 9: Comprehensive Review of Data-Driven Methods for Endovascular Procedure Scene Perception (Continued)
@@
2.2.1 Guidewire Detection and Tracking
@@
Recent research has emphasized clinical integration and real-time performance. Zhang et al. [37] developed a system for real-time intra-operative tracking with low latency (Dice: 0.899), addressing the critical need for seamless integration into clinical workflows. Their subsequent work [80] introduced the Jigsaw transformer framework, incorporating background reverse attention and a novel training strategy to effectively handle multi-wire scenarios (F1: 0.899 for single wire, 0.846 for dual wire). Ravigopal et al. [81] advanced the field with a real-time pose estimation system that combines continuum kinematic modeling with pose optimization, achieving an accuracy of 0.900. This approach underscores the potential for more comprehensive guidewire tracking beyond simple detection.
@@
2.2.2 Catheter Detection and Tracking
@@
2.3 Multi-modal Registration and Integration in Cardiovascular Imaging
@@
Table 10: Summary of Data-Driven Control Methods for Endovascular Intervention Robots
@@
Table 11: Summary of Data-Driven Control Methods for Endovascular Intervention Robots (continued)
No. Authors (Year) Task Category Method Performance Metrics
@@
3 Learning-Based Control for Robotic-Assisted Endovascular Procedures
@@ 4 Challenges
4.1 Underdevelopment of Transformer-Based Models and Foundational Vision Models for Endovascular Procedures
@@
4.2 Scarcity of High-Quality Coronary Imaging Data
@@
4.3 Lack of Unified Workflow Recognition for Endovascular Procedures
@@
4.4 Absence of a Standardized, Large-Scale Benchmark
Recent advancements in large-scale benchmarking, exemplified by the Behavior-1K dataset [149] developed by Fei-Fei Li’s team, have significantly enhanced performance evaluation in endovascular interventions (EIs) by capturing 1,000 diverse, real-world activities within controlled simulation environments. The work deemed ”Best Paper at ICRA 2024” also extended robotic learning through the introduction of standardized, high-quality datasets and models that support reproducibility and generalization across a wide range of robotic systems [150]. In contrast, endovascular robotics remains constrained by the scarcity of large-scale, domain-specific benchmarks that encapsulate the intricacies of human vasculature and the clinical complexity encountered in real-world settings. This limitation hampers systematic comparisons of algorithmic performance, impedes meaningful evaluation under varying clinical conditions, and results in fragmented, non-generalizable findings.
@@ 4.4 Absence of a Standardized, Large-Scale Benchmark
4.5 Scalability of Perception and Learning-Based Control
@@
To address these scalability challenges, future research must concentrate on developing advanced perception algorithms capable of maintaining high accuracy and robustness across a wider range of anatomical variations and pathological conditions. One promising direction involves the integration of multi-modal imaging techniques, including ultrasound and DSA [154], combined with advanced deep learning algorithms, to enhance the system’s capacity to adapt to real-time physiological changes. Additionally, transfer learning and domain adaptation approaches may aid in overcoming the challenge of generalizing control algorithms across different patient datasets. These strategies have the potential to substantially improve the flexibility and clinical applicability of endovascular robotic systems, ensuring their reliable performance across diverse patient populations and ultimately leading to improved procedural outcomes and patient safety.
4.6 Ethical and Legal Concerns
@@
Decision-making autonomy represents another critical issue. While AI-powered robotic systems can enhance precision, their ”closed-box” nature, with opaque decision-making processes, raises concerns about clinician trust [157, 158]. Trustworthiness, defined as the reliability and predictability of a system, is crucial to ensure clinicians have confidence in the technology’s performance. Furthermore, XAI is essential for transparency, enabling clinicians to understand the rationale behind the robot’s decisions. This is particularly important in high-risk procedures, in which clinicians may need to intervene if the system’s decision-making process is unclear [159].
@@
4.7 Integration with Existing Clinical Workflows
Beyond the technical challenges, the integration of robotic systems into existing clinical workflows poses a critical barrier to their widespread adoption. Although the potential of endovascular robotics to enhance precision, decrease procedural variability, and improve patient outcomes is evident, the transition from traditional manual methods to fully automated or semi-automated systems presents substantial challenges beyond purely technological aspects. A pivotal issue is the reluctance of clinical staff to embrace new systems, particularly when they are perceived as ”black boxes” - autonomous systems with decision-making processes that are difficult to interpret or explain [158]. In endovascular procedures, in which precision and safety are paramount, the inability to fully comprehend how a robotic system arrives at its decisions can elicit significant levels of anxiety among medical professionals. Clinicians require assurances that these systems can manage complex and unpredictable situations in a manner akin to experienced human practitioners, and that they would operate predictably in high-pressure clinical environments [163].
@@ 5 Conclusion
Acknowledgement
The authors would like to thank Mr. Mish Toszeghi from Queen Mary University of London for his meticulous proofreading of this manuscript. The authors also appreciate Mr. Haoyu Wang from Shanghai University of Engineering Science for his constructive discussions and valuable suggestions. This work is supported by the National Key Research and Development Program of China under Grant No. 2023YFB4705200, the National Natural Science Foundation of China under Grant No. 62273257, and the Open Project Fund of State Key Laboratory of Cardiovascular Diseases No. 2024SKL-TJ002.
Competing interests
The authors declare no competing interests.